OpenStudy (anonymous):

\[ \sum_{n=1}^{infty} {n!}/{100}^{n}\]

OpenStudy (anonymous):

convergence

OpenStudy (turingtest):

ratio test?

OpenStudy (anonymous):

(n+1)! ------ 100^(n+1) ------ n! ---- 100^n

OpenStudy (turingtest):

\[\lim_{n\to\infty}|\{a_{n+1}\}\cdot\frac1{\{a_n\}}|=\lim_{n\to\infty}|\frac{100^{n+1}}{(n+1)!}\cdot\frac{n!}{100^n}|=\lim_{n\to\infty}|\frac{100}{n+1}|=0<1\]

OpenStudy (turingtest):

I think...

OpenStudy (turingtest):

oh no it's upside-down!

OpenStudy (anonymous):

yeah..but how do we get rid of the "n!"

OpenStudy (turingtest):

\[\lim_{n\to\infty}|\{a_{n+1}\}\cdot\frac1{\{a_n\}}|=\lim_{n\to\infty}|\frac{(n+1)!}{100^{n+1}}\cdot\frac{100^n}{n!}|=\lim_{n\to\infty}|\frac{n+1}{100}|=DNE\]so divergent I guess

OpenStudy (anonymous):

how did you solve the (n+1)! ? I don't know what to do with it

OpenStudy (turingtest):

I just took the limit as n goes to infty\[\lim_{n\to\infty}|\{a_{n+1}\}\cdot\frac1{\{a_n\}}|=\lim_{n\to\infty}|\frac{(n+1)!}{100^{n+1}}\cdot\frac{100^n}{n!}|=\lim_{n\to\infty}|\frac{n+1}{100}|=\frac\infty{100}\]which is a non-existent limit

OpenStudy (anonymous):

Sorry don't mean to be so stubborn. I understand that it's divergent. But here is my issue... on a test, can I just say that lim n-> of (n+1)! = (n+1)

OpenStudy (anonymous):

lim -> infinity of ...

OpenStudy (anonymous):

lim n-> infinity... sorry i can't type today

OpenStudy (anonymous):

i look at the equation and see the "!" sign and don't know what to do with it in term of limits

OpenStudy (anonymous):

Lolz...is my question really that silly?

OpenStudy (turingtest):

sorry I took so long to reply

OpenStudy (anonymous):

no worries

OpenStudy (turingtest):

are you stuck on\[\frac{(n+1)!}{n!}=n+1\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

well look at what happens with factorial with smaller numbers...

OpenStudy (turingtest):

\[\frac{4!}{2!}={4\cdot3\cdot\cancel2\cdot\cancel1\over\cancel2\cdot\cancel1}=4\cdot3=12\]so certain numbers cancel

OpenStudy (anonymous):

got it...so since it's (n+1)!/n! ( n+1) will be the only term standing

OpenStudy (turingtest):

if the number on the bottom is one less than that on the top\[\frac{6!}{5!}={6\cdot\cancel5\cdot\cancel4\cdot\cancel3\cdot\cancel2\cdot\cancel1\over\cancel5\cdot\cancel4\cdot\cancel3\cdot\cancel2\cdot\cancel1}=6\]so yeah, you got the idea

OpenStudy (anonymous):

great! thanks!

OpenStudy (turingtest):

welcome :)